OpenStudy (anonymous):

Given the two sides of a triangle below, determine a possible length of the third side? Select one: A. 28 B. 19.5 C. 28.5 D. 33

OpenStudy (anonymous):



OpenStudy (anonymous):

@quickstudent @mathstudent55

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

use Pythagorean theorem

OpenStudy (anonymous):

i did and got 10.58

OpenStudy (anonymous):

im not sure what i've done wrong... i know im looking for the second leg so i didnt use 16 as b....

Nnesha (nnesha):

yeah right because 16 is hypotheses so it stand for c in the equation

OpenStudy (anonymous):

here is my work.... 12^2+b^2=16^2. 144+b^2=256 b^2=112 square root of 112=10.58.. but that's not an answer choice

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

@study100

OpenStudy (anonymous):

@thelovelyemilyjane Remember Any side of a triangle must be less than the other two sides added together 12+16 = 28 first answer: 28 is equal to 28, so it's not an option, MUST be less than looks like 19.5 is the only option.

OpenStudy (anonymous):

Thank you so much for your help :)

OpenStudy (anonymous):

You're welcomed :)

OpenStudy (anonymous):

would you mind helping me with another one?

OpenStudy (anonymous):

Sure :)

OpenStudy (anonymous):

Of the following degree measures of angle z, which would create the longest length of side JK?

OpenStudy (anonymous):